# CAT: A Conditional Adaptation Tailor for Efficient and Effective Instance-Specific Pansharpening on Real-World Data

### v1

@@ Abstract.
1. Introduction
Remote sensing images are extensively utilized in various domains, such as environmental monitoring (Fitzner et al., 2013; Crampton et al., 2013), urban planning (Pettit et al., 2014; Cheshire and Batty, 2012), and precision agriculture (Zhang and Kovacs, 2012; Mulla, 2013). However, due to hardware limitations, satellite sensors cannot capture images that are both spectrally rich and of high spatial resolution. Typically, satellites like WorldView series capture low-resolution multispectral (LRMS) images and high-resolution panchromatic (PAN) images separately. Consequently, pansharpening aims to fuse an LRMS image with a high-resolution PAN image to generate a high-resolution multispectral (HRMS) image that preserves both spectral and spatial information, which plays a vital role in the field of remote sensing image processing.
Pansharpening methods are broadly classified into four categories: component substitution (CS), multi-resolution analysis (MRA), variational optimization (VO), and deep learning (DL) approaches (Deng et al., 2022; Vivone et al., 2025, 2020). The first three categories are traditional approaches that require only a PAN-LRMS image pair as input. Although CS and MRA methods are computationally efficient, they often suffer from spectral and spatial distortions, respectively. VO approaches rely on the formulation of a variational model. However, they are generally limited in capturing the nonlinear relationships between PAN and HRMS images, which constrains their overall fusion performance (Xiao et al., 2022, 2023). In contrast, deep learning methods leverage neural network architectures excellent feature extraction capability, such as convolutional neural network (CNN)-based models (e.g., (He et al., 2019; Yang et al., 2017; Cao et al., 2021; Yang et al., 2023; Wang et al., 2024)), Transformer-based models (e.g., (Zhou et al., 2022; Li et al., 2024; Zhang and Ma, 2021; Wu et al., 2025)), and Diffusion-based methods (e.g., (Meng et al., 2023a; Cao et al., 2024a; Zhong et al., 2024; Rui et al., 2024b)), enabling them to produce superior fusion results.
@@ 1. Introduction
To address these challenges, we propose an innovative pansharpening framework shown in Figure 1 that incorporates a lightweight, test-time adaptive module, termed the Conditional Adaptation Tailor (CAT). Unlike conventional deep learning approaches that require either time-consuming retraining or computationally intensive zero-shot inference (Rui et al., 2024a), our CAT module is integrated and updated exclusively at test time based solely on the input data, eliminating the need for large-scale retraining. This design seamlessly integrates the CAT module into various pre-trained backbone networks without altering their original parameters, preserving the pretrained model’s performance while enabling rapid adaptation to new data. Specifically, the CAT module is trained in a physical-information-motivated unsupervised manner and is positioned between the backbone’s feature extraction and channel transformation stages, where it learns residual adjustments in the high-dimensional latent representations generated by the early layers. This unsupervised training approach, combined with its strategic placement, allows CAT to efficiently tailor intermediate features to match domain shifts encountered at test time, significantly enhancing cross-dataset and cross-sensor generalization while retaining valuable pretrained knowledge. Furthermore, we employ a patch-wise training and inference strategy: a small number of randomly selected image patches are used to efficiently optimize the CAT parameters, and parallel inference across all patches minimizes overall processing time, enabling high-quality pansharpening of large-scale, megapixel satellite imagery within seconds. With unsupervised learning and conditional feature adaptation, our framework delivers significant gains in both performance and computational efficiency, offering a practical and scalable solution for real-time remote sensing applications. Experiments demonstrate that the proposed method with CAT achieves state-of-the-art performance in real-time application on real-world data.
@@
2. Related Works
2.1. Deep Learning-based Pansharpening
Deep learning has substantially enhanced pansharpening performance by enabling models to capture complex spatial-spectral relationships through various neural network architectures. CNN-based methods such as PanNet (Yang et al., 2017), DiCNN (He et al., 2019), and FusionNet (Deng et al., 2020) have shown strong capabilities in enhancing spatial details while preserving spectral fidelity. Building on this foundation, Transformer-based approaches (Zhou et al., 2022; Li et al., 2024) and diffusion-based models (Meng et al., 2023b; Zhang et al., 2024) further improve performance by leveraging more sophisticated architectures to extract and fuse spatial and spectral features effectively.
@@ 2.1. Deep Learning-based Pansharpening
2.2. Instance-Specific Pansharpening
@@
Traditional pansharpening methods, including CS (Choi et al., 2011; Vivone, 2019a), MRA (Vivone et al., 2014, 2018a), and VO methods (Fu et al., 2019; Tian et al., 2022), only require a single PAN-LRMS pair to generate the fused HRMS image. Efficient as they are, their pansharpening capabilities are often inferior than deep learning-based methods since traditional approaches rely heavily on handcrafted transformations and struggle to capture the complex, nonlinear spatial-spectral relationships inherent in remote sensing data.
On the other hand, zero-shot pansharpening approaches also only require PAN-LRMS pair as input, but also adequately facilitate superior pansharpening capabilities of neural networks. Zero-shot pansharpening frameworks like PSDip (Rui et al., 2024a) and ZS-Pan (Cao et al., 2024b) generally exhibit improved results over traditional approaches in terms of instance-specific pansharpening. Despite efficacy, their processing time towards a single PAN-LRMS pair always prolongs to minute-level, which means their training time is hundreds time longer than inference, making it hardly practical for real-time application.
3. Methodology
@@ 3.1. Overall Framework
3.2. CAT Module Position and Structure Design
@@ 3.2.1. Placement Design
where denotes the fused HRMS inferred by the original backbone network , and denotes the parameters of backbone network. Despite the diverse designs of the backbone network across different approaches, many methods (e.g., (Yang et al., 2017; He et al., 2019; Deng et al., 2020; Liang et al., 2022)) adopt a similar structure that can be easily decomposed into a feature extraction stage and a channel transformation stage. The feature extraction stage is typically implemented by the early layers of the network, denoted as with parameters , while the channel transformation stage is handled by the later layers, denoted as with parameters . Consequently, the overall network can be expressed as , with the complete set of parameters given by .
@@ 3.2.2. Structure Design
3.3. Unsupervised CAT Training Loss
@@ 3.3.1. Spectral Loss ()
3.3.2. Spatial Loss ()
The spatial loss encourages the tailored HRMS image to retain high-frequency spatial details consistent with structural information provided by the PAN image. Similar to the methodology presented in (Wu et al., 2022), the spatial loss is calculated by extracting high-frequency components from both the PAN image and the luminance channel of the generated HRMS image and measuring their difference:
@@ 3.3.2. Spatial Loss ()
3.3.3. Original-Output Consistency Loss ()
@@ 3.3.4. Overall Loss
This unsupervised loss formulation ensures rapid yet effective training, allowing AdaPan to achieve superior performance and excellent generalization across diverse scenarios without incurring significant computational overhead.
3.4. Patch-wise Adaptation and Parallel Inference
@@
4. Experiments
4.1. Datasets, Metrics, and Training Details
@@
Our experiments employ two datasets derived from PanCollection111https://github.com/liangjiandeng/PanCollection, captured by WorldView-3 (WV-3) and WorldView-2 (WV-2), and constructed in accordance with Wald’s protocol (Wald et al., 1997). These datasets consist of panchromatic (PAN) images and low-resolution multispectral (LRMS) images, with PAN images sized at . Notably, our method specifically targets real-world pansharpening challenges—data of paramount practical significance and precisely the type for which our unsupervised loss function was designed. Consequently, we primarily present results on real-world data in the following sections. Additionally, we extend our evaluation to a megapixel image (PAN size: ).
@@
We use well-established evaluation metrics to assess our method quantitatively. For real-world data, we employ HQNR (Aiazzi et al., 2014), and , where HQNR is derived from and to provide a comprehensive assessment of image quality.
@@
Our experimental framework employs FusionNet (Deng et al., 2020) (2020) and U2Net (Peng et al., 2023) (2023) pretrained on WV3 dataset as our baseline backbone networks, representing typical CNN-based and attention-based architectures, respectively. This selection underscores that our method is versatile and can be seamlessly integrated across diverse network designs, enhancing pansharpening performance regardless of the architecture used in the feature extraction layers. And experiments were conducted on a hardware setup comprising an NVIDIA RTX 3090 GPU with 24GB memory and Intel i9-12900 CPU. More details can be found in supplementary materials.
@@
| BT-H (Lolli et al., 2017)2017 | 0.86590.0568 | 0.06560.0262 | 0.07420.0383 | 0.83000.0430 | 0.08600.0301 | 0.09250.0208 |
| C-BDSD (Garzelli, 2014)2014 | 0.85620.0233 | 0.08740.0236 | 0.06180.0128 | 0.69560.0461 | 0.22530.0488 | 0.10190.0296 |
| BDSD-PC (Vivone, 2019b)2019 | 0.86730.0543 | 0.06340.0246 | 0.07490.0359 | 0.82860.0432 | 0.14130.0320 | |
| MTF-GLP (Aiazzi et al., 2006)2006 | 0.90260.0444 | 0.03730.0124 | 0.06280.0359 | 0.85490.0475 | 0.05820.0221 | 0.09300.0320 |
| MTF-GLP-FS (Vivone et al., 2018b)2018 | 0.91270.0348 | 0.03570.0106 | 0.05370.0273 | 0.86580.0415 | 0.05630.0212 | 0.08300.0260 |
| MF (Restaino et al., 2016)2016 | 0.90140.0358 | 0.04520.0121 | 0.05610.0268 | 0.85080.0538 | 0.07040.0308 | 0.08570.0297 |
| PsDip (Rui et al., 2024a)2024 | 0.92150.0176 | 0.01910.0078 | 0.06070.0117 | 0.89800.0226 | 0.06590.0158 | |
| ZS-Pan (Cao et al., 2024b)2024 | 0.94490.0208 | 0.02540.0071 | 0.03060.0153 | 0.04760.0270 | 0.04350.0130 | |
| FusionMamba (Peng et al., 2024)2024 | 0.01860.0078 | 0.90640.0227 | 0.05270.0252 | 0.04310.0114 | ||
| WFANet (Huang et al., 2025)2025 | 0.94630.0095 | 0.03810.0042 | 0.89220.0223 | 0.06330.0221 | 0.04760.0073 | |
| FusionNet (Deng et al., 2020)2020 | 0.94050.0197 | 0.02290.0080 | 0.03750.0143 | 0.88810.0213 | 0.05430.0273 | 0.06060.0162 |
| U2Net (Peng et al., 2023)2023 | 0.95140.0114 | 0.03130.0074 | 0.87060.0809 | 0.09360.0790 | 0.04000.0098 | |
@@
4.2. Benchmark
To demonstrate the effectiveness of our proposed method, we compared it with several instance-specific pansharpening approaches, including BT-H (Lolli et al., 2017), C-BDSD (Garzelli, 2014), BDSD-PC (Vivone, 2019b), MTF-GLP (Aiazzi et al., 2006), MTF-GLP-FS (Vivone et al., 2018b), and MF (Restaino et al., 2016). We also evaluated our method against state-of-the-art zero-shot pansharpening approaches, such as PsDip (Rui et al., 2024a) and ZS-Pan (Cao et al., 2024b). Additionally, we included baseline models—namely, FusionNet (Deng et al., 2020) and U2Net (Peng et al., 2023)—to further validate the performance improvements achieved by our approach.
4.3. Main Experimental Results
@@
4.4. Effectiveness and Generalization Ability
@@ 4.5. Discussions
To validate our method on megapixel-level data, we conducted experiments using a PAN image and a LRMS image as inputs to generate a HRMS image e, comprising a total of 128 million pixels. We then compared the results of our method with six classical instance-specific pansharpening methods.
@@
| BT-H (Lolli et al., 2017)2017 | 0.7098 | 0.1710 | 0.1444 | |
| C-BDSD (Garzelli, 2014)2014 | 0.7475 | 0.2019 | 0.0704 | 104.16 |
| BDSD-PC (Vivone, 2019b)2019 | 0.7369 | 0.1723 | 0.1093 | 4.30 |
| MTF-GLP (Aiazzi et al., 2006)2006 | 0.7606 | 0.1231 | 0.1330 | 13.08 |
| MTF-GLP-FS (Vivone et al., 2018b)2018 | 0.7680 | 0.1272 | 0.1199 | 9.91 |
| MF (Restaino et al., 2016)2016 | 0.7709 | 0.1176 | 0.1268 | 7.66 |
@@
4.6. Limitations and Future Work
@@ Appendix A Training Details
For optimization, we employed the Adam optimizer (Kingma and Ba, 2017), a robust choice for training deep learning models, with the learning rate set to and weight decay at to prevent overfitting. The learning rate was carefully tuned to balance convergence speed with stability. We applied the L1 loss criterion for all components of the loss function, which promotes sparsity in the learned features and enhances generalization, particularly when working with high-dimensional data like remote sensing images.
@@ Appendix B Detailed Discussion
B.1. Influence of Number of Selected Patches
@@
B.2. Influence of the Patch Size

### v2

@@ Abstract
Introduction
Remote sensing images are extensively utilized in various domains, such as environmental monitoring (Fitzner et al. 2013; Crampton et al. 2013), urban planning (Pettit et al. 2014; Cheshire and Batty 2012), and precision agriculture (Zhang and Kovacs 2012; Mulla 2013). However, due to hardware limitations, satellite sensors cannot capture images that are both spectrally rich and of high spatial resolution. Typically, satellites like WorldView series capture low-resolution multispectral (LRMS) images and high-resolution panchromatic (PAN) images separately. Consequently, pansharpening aims to fuse an LRMS image with a high-resolution PAN image to generate a high-resolution multispectral (HRMS) image that preserves both spectral and spatial information, which plays a vital role in the field of remote sensing image processing.
Pansharpening methods are broadly classified into four categories: component substitution (CS), multi-resolution analysis (MRA), variational optimization (VO), and deep learning (DL) approaches (Deng et al. 2022; Vivone et al. 2025, 2020). The first three categories are traditional approaches that require only a PAN-LRMS image pair as input. Although CS and MRA methods are computationally efficient, they often suffer from spectral and spatial distortions, respectively. VO approaches rely on the formulation of a variational model. However, they are generally limited in capturing the nonlinear relationships between PAN and HRMS images, which constrains their overall fusion performance (Xiao et al. 2022, 2023). In contrast, deep learning methods leverage neural network architectures excellent feature extraction capability, such as convolutional neural network (CNN)-based models (e.g., (He et al. 2019; Yang et al. 2017; Cao et al. 2021; Yang et al. 2023; Wang et al. 2024)), Transformer-based models (e.g., (Zhou, Liu, and Wang 2022; Li et al. 2024; Zhang and Ma 2021; Wu et al. 2025)), and Diffusion-based methods (e.g., (Meng et al. 2023a; Cao et al. 2024b; Zhong et al. 2024; Rui et al. 2024b)), enabling them to produce superior fusion results.
@@ Introduction
To address these challenges, we propose an innovative pansharpening framework shown in Figure 1 that incorporates a lightweight, test-time adaptive module, termed the Conditional Adaptation Tailor (CAT). Unlike conventional deep learning approaches that require either time-consuming retraining or computationally intensive zero-shot inference (Rui et al. 2024a), our CAT module is integrated and updated exclusively at test time based solely on the input data, eliminating the need for large-scale retraining. This design seamlessly integrates the CAT module into various pre-trained backbone networks without altering their original parameters, preserving the pretrained model’s performance while enabling rapid adaptation to new data. Specifically, the CAT module is trained in a physical-information-motivated unsupervised manner and is positioned between the backbone’s feature extraction and channel transformation stages, where it learns residual adjustments in the high-dimensional latent representations generated by the early layers. This unsupervised training approach, combined with its strategic placement, allows CAT to efficiently tailor intermediate features to match domain shifts encountered at test time, significantly enhancing cross-dataset and cross-sensor generalization while retaining valuable pretrained knowledge. Furthermore, we employ a patch-wise training and inference strategy: a small number of randomly selected image patches are used to efficiently optimize the CAT parameters, and parallel inference across all patches minimizes overall processing time, enabling high-quality pansharpening of large-scale, megapixel satellite imagery within seconds. With unsupervised learning and conditional feature adaptation, our framework delivers significant gains in both performance and computational efficiency, offering a practical and scalable solution for real-time remote sensing applications. Experiments demonstrate that the proposed method with CAT achieves state-of-the-art performance in real-time application on real-world data.
@@
Related Works
Deep Learning-based Pansharpening
Deep learning has substantially enhanced pansharpening performance by enabling models to capture complex spatial-spectral relationships through various neural network architectures. CNN-based methods such as PanNet (Yang et al. 2017), DiCNN (He et al. 2019), and FusionNet (Deng et al. 2020) have shown strong capabilities in enhancing spatial details while preserving spectral fidelity. Building on this foundation, Transformer-based approaches (Zhou, Liu, and Wang 2022; Li et al. 2024) and diffusion-based models (Meng et al. 2023b; Zhang et al. 2024) further improve performance by leveraging more sophisticated architectures to extract and fuse spatial and spectral features effectively.
@@
Instance-Specific Pansharpening
@@
Traditional pansharpening methods, including CS (Choi, Yu, and Kim 2011; Vivone 2019a), MRA (Vivone et al. 2014; Vivone, Restaino, and Chanussot 2018a), and VO methods (Fu et al. 2019; Tian et al. 2022), only require a single PAN-LRMS pair to generate the fused HRMS image. Efficient as they are, their pansharpening capabilities are often inferior than deep learning-based methods since traditional approaches rely heavily on handcrafted transformations and struggle to capture the complex, nonlinear spatial-spectral relationships inherent in remote sensing data.
On the other hand, zero-shot pansharpening approaches also only require PAN-LRMS pair as input, but also adequately facilitate superior pansharpening capabilities of neural networks. Zero-shot pansharpening frameworks like PSDip (Rui et al. 2024a) and ZS-Pan (Cao et al. 2024a) generally exhibit improved results over traditional approaches in terms of instance-specific pansharpening. Despite efficacy, their processing time towards a single PAN-LRMS pair always prolongs to minute-level, which means their training time is hundreds time longer than inference, making it hardly practical for real-time application.
Methodology
@@
CAT Module Position and Structure Design
@@
where denotes the fused HRMS inferred by the original backbone network , and denotes the parameters of backbone network. Despite the diverse designs of the backbone network across different approaches, many methods (e.g., (Yang et al. 2017; He et al. 2019; Deng et al. 2020; Liang et al. 2022)) adopt a similar structure that can be easily decomposed into a feature extraction stage and a channel transformation stage. The feature extraction stage is typically implemented by the early layers of the network, denoted as with parameters , while the channel transformation stage is handled by the later layers, denoted as with parameters . Consequently, the overall network can be expressed as , with the complete set of parameters given by .
@@
Unsupervised CAT Training Loss
@@
Spatial Loss ()
The spatial loss encourages the tailored HRMS image to retain high-frequency spatial details consistent with structural information provided by the PAN image. Similar to the methodology presented in (Wu et al. 2022), the spatial loss is calculated by extracting high-frequency components from both the PAN image and the luminance channel of the generated HRMS image and measuring their difference:
@@
Original-Output Consistency Loss ()
@@
This unsupervised loss formulation ensures rapid yet effective training, allowing the CAT module to achieve superior performance and excellent generalization across diverse scenarios without incurring significant computational overhead.
Patch-wise Adaptation and Parallel Inference
@@
Experiments
Datasets, Metrics, and Training Details
@@
Our experiments employ two datasets derived from PanCollection111https://github.com/liangjiandeng/PanCollection, captured by WorldView-3 (WV-3) and WorldView-2 (WV-2), and constructed in accordance with Wald’s protocol (Wald, Ranchin, and Mangolini 1997). These datasets consist of panchromatic (PAN) images and low-resolution multispectral (LRMS) images, with PAN images sized at . Notably, our method specifically targets real-world pansharpening challenges—data of paramount practical significance and precisely the type for which our unsupervised loss function was designed. Consequently, we primarily present results on real-world data in the following sections. Additionally, we extend our evaluation to a megapixel image (PAN size: ).
@@
We use well-established evaluation metrics to assess our method quantitatively. For real-world data, we employ HQNR (Aiazzi et al. 2014), and , where HQNR is derived from and to provide a comprehensive assessment of image quality.
@@
Our experimental framework employs FusionNet (Deng et al. 2020) (2020) and U2Net (Peng et al. 2023) (2023) pretrained on WV3 dataset as our baseline backbone networks, representing typical CNN-based and attention-based architectures, respectively. This selection underscores that our method is versatile and can be seamlessly integrated across diverse network designs, enhancing pansharpening performance regardless of the architecture used in the feature extraction layers. And experiments were conducted on a hardware setup comprising an NVIDIA RTX 3090 GPU with 24GB memory and Intel i9-12900 CPU. More details can be found in supplementary materials.
@@
| BT-H 2017 | 0.86590.0568 | 0.06560.0262 | 0.07420.0383 | 0.83000.0430 | 0.08600.0301 | 0.09250.0208 |
| C-BDSD 2014 | 0.85620.0233 | 0.08740.0236 | 0.06180.0128 | 0.69560.0461 | 0.22530.0488 | 0.10190.0296 |
| BDSD-PC2019 | 0.86730.0543 | 0.06340.0246 | 0.07490.0359 | 0.82860.0432 | 0.14130.0320 | |
| MTF-GLP 2006 | 0.90260.0444 | 0.03730.0124 | 0.06280.0359 | 0.85490.0475 | 0.05820.0221 | 0.09300.0320 |
| MTF-GLP-FS 2018 | 0.91270.0348 | 0.03570.0106 | 0.05370.0273 | 0.86580.0415 | 0.05630.0212 | 0.08300.0260 |
| MF 2016 | 0.90140.0358 | 0.04520.0121 | 0.05610.0268 | 0.85080.0538 | 0.07040.0308 | 0.08570.0297 |
| PsDip 2024 | 0.92150.0176 | 0.01910.0078 | 0.06070.0117 | 0.89800.0226 | 0.06590.0158 | |
| ZS-Pan 2024 | 0.94490.0208 | 0.02540.0071 | 0.03060.0153 | 0.04760.0270 | 0.04350.0130 | |
| FusionMamba 2024 | 0.01860.0078 | 0.90640.0227 | 0.05270.0252 | 0.04310.0114 | ||
| WFANet 2025 | 0.94630.0095 | 0.03810.0042 | 0.89220.0223 | 0.06330.0221 | 0.04760.0073 | |
| FusionNet 2020 | 0.94050.0197 | 0.02290.0080 | 0.03750.0143 | 0.88810.0213 | 0.05430.0273 | 0.06060.0162 |
| U2Net 2023 | 0.95140.0114 | 0.03130.0074 | 0.87060.0809 | 0.09360.0790 | 0.04000.0098 | |
@@
Benchmark
To demonstrate the effectiveness of our proposed method, we compared it with several instance-specific pansharpening approaches, including BT-H (Lolli et al. 2017), C-BDSD (Garzelli 2014), BDSD-PC (Vivone 2019b), MTF-GLP (Aiazzi et al. 2006), MTF-GLP-FS (Vivone, Restaino, and Chanussot 2018b), and MF (Restaino et al. 2016). We also evaluated our method against state-of-the-art zero-shot pansharpening approaches, such as PsDip (Rui et al. 2024a) and ZS-Pan (Cao et al. 2024a). Additionally, we included baseline models—namely, FusionNet (Deng et al. 2020) and U2Net (Peng et al. 2023)—to further validate the performance improvements achieved by our approach.
Main Experimental Results
@@
Effectiveness and Generalization Ability
@@
To validate our method on megapixel-level data, we conducted experiments using a PAN image and a LRMS image as inputs to generate a HRMS image, comprising a total of 128 million pixels. We then compared the results of our method with six classical instance-specific pansharpening methods.
@@
| BT-H 2017 | 0.7098 | 0.1710 | 0.1444 | |
| C-BDSD 2014 | 0.7475 | 0.2019 | 0.0704 | 104.16 |
| BDSD-PC2019 | 0.7369 | 0.1723 | 0.1093 | 4.30 |
| MTF-GLP 2006 | 0.7606 | 0.1231 | 0.1330 | 13.08 |
| MTF-GLP-FS 2018 | 0.7680 | 0.1272 | 0.1199 | 9.91 |
| MF 2016 | 0.7709 | 0.1176 | 0.1268 | 7.66 |
@@
Limitations and Future Work
@@ Appendix A Training Details
For optimization, we employed the Adam optimizer (Kingma and Ba 2017), a robust choice for training deep learning models, with the learning rate set to and weight decay at to prevent overfitting. The learning rate was carefully tuned to balance convergence speed with stability. We applied the L1 loss criterion for all components of the loss function, which promotes sparsity in the learned features and enhances generalization, particularly when working with high-dimensional data like remote sensing images.
@@ Appendix B Detailed Discussion
Influence of Number of Selected Patches
@@
Influence of the Patch Size